# Return of the Lepton Number: Sterile Neutrino Dark Matter Production and the Revival of the Shi-Fuller Mechanism

### v1

@@ II.2 Lepton Number Evolution
In order to best utilize recent findings, we will treat the lepton asymmetry as being unconstrained at the order $L\sim 10$, based on Froustey and Pitrou (FP) [54] and in contrast to previous treatments of SF production that considered $L$ as evolving through a synchronized neutrino oscillation equalization process, and thus placed a constraint on models utilizing $L>0.05$ [61, 62, 63]. FP revisit cosmological constraints on primordial lepton asymmetries using a full three-flavor, multi-momentum neutrino transport code that solves the quantum kinetic equations with exact collision terms and self-consistently couples to a Big Bang Nucleosynthesis (BBN) solver [54]. The authors demonstrate that in order to satisfy BBN bounds, one need not have asymmetries of order $L\sim 0.1$, or smaller, but that in scenarios with large or flavor-imbalanced initial asymmetries, initially very large asymmetries at $T\gg 1\,\mathrm{MeV}$ can equilibrate to small, unconstrained $L_{\alpha}$ when entering the BBN era. We illustrate these results specifically for the parameter space we are interested in in Section II.2.2.
Finally, we note that the lepton potential in sterile neutrino dark matter production depends on the total lepton number in an asymmetric way, with a factor of 2 greater dependence on the lepton number in the species with which the sterile neutrino mixes. The lepton potential is [24]
@@ II.2.1 Cosmic Lepton Number Evolution Above $T\approx 20\rm\,MeV$
Our calculations start with asymmetries well-prior to all dark matter production at $T=10\,\mathrm{GeV}$, and end after all dark matter production, at $T=10\,\mathrm{MeV}$. It was found in initial work [61, 62, 63] and in FP that large asymmetries are largely frozen in evolution at $T\gg 10\,\mathrm{MeV}$ due to any active-neutrino mixing. However, during sterile neutrino dark matter production, $L$ is affected by conversion to the sterile neutrinos’ “lepton number,” redistribution of any initial neutrino asymmetry into charged lepton and hadronic asymmetries, as well as redistribution into the other neutrino flavor asymmetries [57].
The redistribution of lepton number during dark matter production is computed self-consistently in sterile-dm by solving a set of coupled linear equations that enforce the chemical and kinetic equilibrium among the various components in the plasma [57]. These constraint equations arise from the relation between the lepton number asymmetries and the corresponding chemical potential and number-density susceptibility of each lepton (both charged and neutral), as well as relations describing charge and baryon number conservation. The temperature dependent susceptibilities are carefully constructed across three epochs: using perturbative QCD at temperatures well above the quark-hadron transition ($T\gtrsim 300\,\mathrm{MeV}$), lattice QCD near the transition ($150\,\mathrm{MeV}\lesssim T\lesssim 300\,\mathrm{MeV}$), and a hadron resonance gas model well below the transition ($T\lesssim 150\,\mathrm{MeV}$), with care taken to smoothly match each regime. This framework captures how the initial neutrino lepton asymmetry in one flavor is dynamically redistributed into both charged lepton asymmetries, as well as into the hadronic sector.
As an example, $\nu_{\mu}+\mu^{+}\rightleftharpoons\pi^{+}$, which is faster than the Hubble rate in the dark matter production period, is one of the primary channels of this redistribution at temperatures near peak production. Redistribution between leptonic flavors becomes more dominant at lower temperatures below the quark-hadron transition. The result of redistribution across the dark matter production period is a $\simeq 10\%$ correction in the asymmetry potential [57]. However, as shown below, the corresponding change in $L$ across the same period is quite large, up to a factor of 10 in the examined region of parameter space.
In Fig. 2, we show an example case of the evolution of $L$ for the successful production of the dark matter for a sterile neutrino with a particle mass of 40 keV and a mixing angle of $\sin^{2}{(2\theta)}=1.6\times 10^{-15}$, corresponding to point B in Fig. 1. As can be seen in the figure, production conversion and redistribution of $L$ leads to its reduction in magnitude by a factor of greater than five.
@@ II.2.2 Cosmic Lepton Number Evolution Below $T\approx 20\rm\,MeV$ and Cosmological Constraints
Below $T\sim 20\,\mathrm{MeV}$, synchronized neutrino oscillations redistribute the asymmetry between the electron, muon, and tau neutrino flavors [61, 62, 63]. Although the general trend is an approximate equilibration of the individual lepton numbers $L_{\alpha}$ to their average $L_{\mathrm{av}}=(\sum_{\alpha}L_{\alpha})/3$, recent studies [64, 54, 55] have shown that the use of modern values of the mixing parameters and the full collision term significantly affect the details of neutrino asymmetry evolution. Lepton asymmetries can be constrained via their effect on cosmological observables, namely, the effective number of relativistic species $N_{\mathrm{eff}}$ and the primordial abundances of helium-4 $Y_{p}$ and deuterium $\mathrm{D/H}$ (see, e.g., Refs. [65, 66, 67]). Given the current experimental data on these various quantities, the strongest bounds come from BBN spectroscopic measurements (at the percent level), which tightly constrain the electron neutrino asymmetry $L_{e}$ at the BBN epoch. Although it had been recognized before that specific configurations, with potentially large individual $L_{\alpha}$, could lead to small values of the final $L_{e}$ (see e.g., [68]), FP have shown that a large domain of the parameter space with $L\gtrsim 1$ at $T\approx 25\,\mathrm{MeV}$ is allowed by experimental data [54].
@@
Figure 3: Shown here are outputs of the numerical solution of the neutrino quantum kinetic equations and self-consistent BBN calculations for initial asymmetry configurations at $T=20\,\mathrm{MeV}$ with $L_{\mu}\approx 0.68$, varying $L_{e}$ and the average lepton number $L_{\mathrm{av}}$. Left panels show helium-4 (top) and deuterium (bottom) abundances for a baryon density $\Omega_{b}h^{2}=0.0224$ (preferred value from Planck [69]). The dotted (dashed) lines are the 68 % (95 %) confidence levels from the spectroscopic measurements of $Y_{p}$ [70] and $\mathrm{D/H}$ [71, 72, 73]. Top right panel shows the effective number of relativistic species $N_{\mathrm{eff}}$, which is always larger than the standard value $3.044$ at zero asymmetries [74, 75, 76]. Bottom right panel is the likelihood based on CMB and BBN measurements (see details in [54]), compared to the likelihood ~$\tilde{\mathcal{L}}_{0}$ of the symmetric baseline $L_{e,\mu,\tau}=0$.
@@
In what follows, we show specifically that many asymmetry configurations with a large $L_{\mu}$ (which is the case considered here for dark matter production) are compatible with experimental constraints. For that purpose, we apply the numerical code used in Ref. [54] (see also Ref. [64]) to a range of asymmetry configurations at $T=20\,\mathrm{MeV}$, with a fixed muon neutrino asymmetry $\eta_{\mu}=1/6$, that is, $L_{\mu}\approx 0.68$. We assume a normal ordering of neutrino masses and use the mixing parameters from [77]. We show in Fig. 3 the various cosmological observables ($Y_{p}$, $\mathrm{D/H}$ and $N_{\mathrm{eff}}$) obtained for each initial configuration. For primordial nucleosynthesis, the (anti)neutrino distributions obtained from the solution of the quantum kinetic equations are used as input in the BBN code PRIMAT [78]; we show the resulting abundances obtained for the preferred value of the baryon density from Planck ($\Omega_{b}h^{2}=0.0224$ [69]). On the bottom right panel, we compute the likelihood $\tilde{\mathcal{L}}$ of each point, as a combination of gaussian likelihoods for each BBN spectroscopic measurement (helium-4 [70], deuterium [71, 72, 73]) and the likelihood associated to the combination of CMB data described in Ref. [54]. We then plot the ratio of this likelihood to that of the lepton-symmetric configuration $(L_{e},L_{\mu},L_{\tau})=(0,0,0)$. For reference, the point of highest likelihood in this limited region of the parameter space is actually preferred to the “standard model” origin, with $-2\ln(\tilde{\mathcal{L}}_{\mathrm{max}}/\tilde{\mathcal{L}}_{0})\approx-0.94$. For visualization purposes, we represent the confidence levels of the spectroscopic measurements of primordial abundances on the left panels of Fig. 3. The shift between the preferred regions for $Y_{p}$ and $\mathrm{D/H}$ stems from the “deuterium tension” in PRIMAT [79], which will not be resolved or confirmed until better determinations of the deuterium burning nuclear rates are available [80].
@@
Figure 4 shows the evolution of neutrino asymmetries below $T=20\,\mathrm{MeV}$ for two configurations: on the top panel, a preferred point from the parameter space explored in Fig. 3 (shown with a black star), and on the bottom panel a configuration outside of the range of Fig. 3. Since the temperature range shown on Fig. 4 covers neutrino decoupling and electron-positron annihilations, we need to distinguish between the photon temperature $T_{\gamma}$ and the comoving temperature $T_{\mathrm{cm}}\propto a^{-1}$ (with $a$ the scale factor), which would be the neutrino temperature in the instantaneous neutrino decoupling approximation. Because of the variation of $T_{\gamma}/T_{\mathrm{cm}}$, the lepton number $L_{\alpha}$ defined in Eq. (3) is not a comoving quantity, which is why we multiply it by $(T_{\gamma}/T_{\mathrm{cm}})^{3}$ in Fig. 4 (see also discussion in Ref [54]).
@@
These two examples result in a very small electron neutrino asymmetry at the BBN epoch, which is the main parameter determining BBN abundances. We note that future Simons Observatory and CMB-S4-like observations will be able to constrain $N_{\mathrm{eff}}$ at much better levels, and measure $Y_{p}$ at a level of precision comparable to current spectroscopic measurements [81, 82, 83], thus being able to identify or constrain these required $L$ for the SF mechanism.
@@
With the results of Figs. 3 and 4, we highlight the variety of currently allowed asymmetry configurations for a given muon lepton number $L_{\mu}$. In the following, we then consider the lepton number $L$ which enters dark matter production as being unconstrained at the order $L\lesssim 10$ (recalling that it decreases by a factor of at least 5 between $T=10\rm\,GeV$ and $T=20\,\mathrm{MeV}$).
@@ II.3 Cosmological Structure
WDM affects cosmological structure formation due to free-streaming of the dark matter. The nature of the free-streaming is unique in the case of thermal WDM, but widely varied in the case of non-thermal WDM such as DW and SF sterile neutrino dark matter. The effects of free-streaming are imprinted on the distribution of matter and affect the linear matter power spectrum, and related nonlinear cosmological observables. To produce the linear matter power spectra for each generated SF PSD, we used the CLASS Boltzmann code [84]. The benchmark parameters for $\Lambda$CDM are from the Planck 2018 minimal case of one massive neutrino species with $\Sigma m_{\nu}=0.06$ eV, modeled as a non-CDM (NCDM) dark matter component [69]. Specifically, we adopt the cosmological parameters that are the best estimate parameters from the TT,TE,EE+lowE+lensing column in Table 2 of Ref. [69]. To produce the matter power spectra for the SF cases, we provide the PSDs and particle masses to CLASS. CLASS rescales the PSDs to match its normalization conventions.
We make the assumption that all lepton asymmetries are small by the time of structure formation. That is, they can correspond to lepton number configurations such as in the top panel of Fig. 4, so that the asymmetries are small enough that the structure formation is mostly sensitive to the sterile neutrino dark matter component, and not the small remaining active asymmetries. However, in general, SF mechanisms will likely be connected with non-trivial signals in $L_{\alpha}$ and $N_{\mathrm{eff}}$ in the BBN and CMB eras.
@@
and an example is shown in Figure 5. We fit the SF transfer functions with thermal WDM transfer functions, where the thermal WDM particle mass is the free parameter for each sterile neutrino particle mass and mixing angle point. As shown, the residuals between the closest-fit thermal WDM transfer functions and those of the SF-produced sterile neutrinos’ are relatively small ($\lesssim 20\%$) for the suppression scale of 1/2. The $\lesssim 20\%$ accuracy holds in our fit across the examined parameter space, despite many points having PSDs that are highly non-thermal. In order to properly convert most of the prior literature’s inferred thermal WDM particle masses, $m_{\mathrm{th,old}}$ [85] (more accurate at $m_{\mathrm{th}}\equiv m_{\mathrm{th,new}}\lesssim 5\,\mathrm{keV}$) into that calculated for larger thermal WDM particle masses, $m_{\mathrm{th,new}}$, found in Ref. [21], we use a linear fit relation,
@@
where the particle masses are in units of keV. This relation should be used to properly map any work which utilizes the previous WDM transfer function fit, in order to convert it into a more broadly calibrated thermal WDM particle mass value. We repeat our fitting process across the full parameter space and derive contours to describe the thermal WDM equivalent masses in order to apply signals and constraints directly to the SF parameter space. This is shown in Section III and Figure 6.
Figure 6: Shown is the full sterile neutrino dark matter parameter space under SF production, with the initial $L$ values given at $T=10\rm\,GeV$ in order to achieve the observed density of the dark matter. These $L$ values are all reduced by a factor of 5 to 10 by $T=10\,\rm MeV$ from dark matter production and dilution (see Sec. II.2.1 and Fig. 2). The $L$ is further reduced by synchronized oscillations prior to BBN, and the initial $L$ values shown are consistent with BBN and CMB observations today (see Sec. II.2.2). The blue dot-dashed line is the best fit the thermal WDM particle mass from Villasenor+ [20]. The full region shown is consistent with the $1\sigma$ thermal WDM particle mass region from Villasenor+. The dashed red contour is the thermal WDM particle mass lower bound from the Zelko+ [19], disfavoring the region left of the line. The gray cross-hatched X-ray exclusion region is the combination of NuSTAR and SPI/INTEGRAL at 95% C.L. upper limits [39, 40, 41, 42, 43].

### v3

@@ II.2 Lepton number evolution
In order to best utilize recent findings, we will treat the lepton asymmetry as being unconstrained at the order $L\sim 10$, based on Froustey and Pitrou (FP) [Froustey:2024mgf] and in contrast to previous treatments of SF production that considered $L$ as evolving through a synchronized neutrino oscillation equalization process, and thus placed a constraint on models utilizing $L>0.05$ [Dolgov:2002ab, Abazajian:2002qx, Wong:2002fa]. FP revisit cosmological constraints on primordial lepton asymmetries using a full three-flavor, multimomentum neutrino transport code that solves the quantum kinetic equations with exact collision terms and self-consistently couples to a big bang nucleosynthesis (BBN) solver [Froustey:2024mgf]. The authors demonstrate that in order to satisfy BBN bounds, one need not have asymmetries of order $L\sim 0.1$, or smaller, but that in scenarios with large or flavor-imbalanced initial asymmetries, initially very large asymmetries at $T\gg 1\,\mathrm{MeV}$ can equilibrate to small, unconstrained $L_{\alpha}$ when entering the BBN era. We illustrate these results specifically for the parameter space we are interested in in Sec. II.2.2.
Finally, we note that the lepton potential in sterile neutrino dark matter production depends on the total lepton number in an asymmetric way, with a factor of 2 greater dependence on the lepton number in the species with which the sterile neutrino mixes. The lepton potential is [Abazajian:2001nj]
@@ II.2.1 Cosmic lepton number evolution above $T\approx 20\rm\,MeV$
Our calculations start with asymmetries well-prior to all dark matter production at $T=10\,\mathrm{GeV}$, and end after all dark matter production, at $T=10\,\mathrm{MeV}$. It was found in initial work [Dolgov:2002ab, Abazajian:2002qx, Wong:2002fa] and in FP that large asymmetries are largely frozen in evolution at $T\gg 10\,\mathrm{MeV}$ due to any active-neutrino mixing. However, during sterile neutrino dark matter production, $L$ is affected by conversion to the sterile neutrinos’ “lepton number,” redistribution of any initial neutrino asymmetry into charged lepton and hadronic asymmetries, as well as redistribution into the other neutrino flavor asymmetries [Venumadhav:2015pla].
The redistribution of the lepton number during dark matter production is computed self-consistently in sterile-dm by solving a set of coupled linear equations that enforce the chemical and kinetic equilibrium among the various components in the plasma [Venumadhav:2015pla]. These constraint equations arise from the relation between the lepton number asymmetries and the corresponding chemical potential and number-density susceptibility of each lepton (both charged and neutral), as well as relations describing charge and baryon number conservation. The temperature dependent susceptibilities are carefully constructed across three epochs: using perturbative QCD at temperatures well above the quark-hadron transition ($T\gtrsim 300\,\mathrm{MeV}$), lattice QCD near the transition ($150\,\mathrm{MeV}\lesssim T\lesssim 300\,\mathrm{MeV}$), and a hadron resonance gas model well below the transition ($T\lesssim 150\,\mathrm{MeV}$), with care taken to smoothly match each regime. This framework captures how the initial neutrino lepton asymmetry in one flavor is dynamically redistributed into both charged lepton asymmetries, as well as into the hadronic sector.
As an example, $\nu_{\mu}+\mu^{+}\rightleftharpoons\pi^{+}$, which is faster than the Hubble rate in the dark matter production period, is one of the primary channels of this redistribution at temperatures near peak production. Redistribution between leptonic flavors becomes more dominant at lower temperatures below the quark-hadron transition. The result of redistribution across the dark matter production period is a $\simeq 10\%$ correction in the asymmetry potential [Venumadhav:2015pla]. However, as shown below, the corresponding change in $L$ across the same period is quite large, up to a factor of 10 in the examined region of parameter space.
In Fig. 3, we show an example case of the evolution of $L$ for the successful production of the dark matter for a sterile neutrino with a particle mass of 40 keV and a mixing angle of $\sin^{2}{(2\theta)}=1.6\times 10^{-15}$, corresponding to point B in Fig. 1. As can be seen in the figure, production conversion and redistribution of $L$ leads to its reduction in magnitude by a factor of greater than 5.
@@ II.2.2 Cosmic lepton number evolution below $T\approx 20\rm\,MeV$ and cosmological constraints
[…]
@@
Figure 4: Shown here are outputs of the numerical solution of the neutrino quantum kinetic equations and self-consistent BBN calculations for initial asymmetry configurations at $T=20\,\mathrm{MeV}$ with $L_{\mu}\approx 0.68$, varying $L_{e}$ and the average lepton number $L_{\mathrm{av}}$. Left panels show helium-4 (top) and deuterium (bottom) abundances for a baryon density $\Omega_{b}h^{2}=0.0224$ (preferred value from Planck [Planck:2018vyg]). The dotted (dashed) lines are the 68 % (95 %) confidence levels from the spectroscopic measurements of $Y_{p}$ [Aver:2020fon] and $\mathrm{D/H}$ [Cooke:2017cwo, Kislitsyn:2024jvk, Guarneri:2024qxi]. Top right panel shows the effective number of relativistic species $N_{\mathrm{eff}}$, which is always larger than the standard value $3.044$ at zero asymmetries [Akita:2020szl, Froustey:2020mcq, Bennett:2020zkv, Drewes:2024wbw]. Bottom right panel is the likelihood based on CMB and BBN measurements (see details in [Froustey:2024mgf]), compared to the likelihood $\tilde{\mathcal{L}}_{0}$ of the symmetric baseline $L_{e,\mu,\tau}=0$.
@@
In what follows, we show specifically that many asymmetry configurations with a large $L_{\mu}$ (which is the case considered here for dark matter production) are compatible with experimental constraints. For that purpose, we apply the numerical code used in Ref. [Froustey:2024mgf] (see also Ref. [Froustey:2021azz]) to a range of asymmetry configurations at $T=20\,\mathrm{MeV}$, with a fixed muon neutrino asymmetry $\eta_{\mu}=1/6$, that is, $L_{\mu}\approx 0.68$. We assume a normal ordering of neutrino masses and use the mixing parameters from [ParticleDataGroup:2024cfk]. We show in Fig. 4 the various cosmological observables ($Y_{p}$, $\mathrm{D/H}$ and $N_{\mathrm{eff}}$) obtained for each initial configuration. For primordial nucleosynthesis, the (anti)neutrino distributions obtained from the solution of the quantum kinetic equations are used as input in the BBN code PRIMAT [Pitrou:2018cgg]; we show the resulting abundances obtained for the preferred value of the baryon density from Planck ($\Omega_{b}h^{2}=0.0224$ [Planck:2018vyg]). On the bottom right panel, we compute the likelihood $\tilde{\mathcal{L}}$ of each point, as a combination of Gaussian likelihoods for each BBN spectroscopic measurement (helium-4 [Aver:2020fon], deuterium [Cooke:2017cwo, Kislitsyn:2024jvk, Guarneri:2024qxi]) and the likelihood associated to the combination of CMB data described in Ref. [Froustey:2024mgf]. We then plot the ratio of this likelihood to that of the lepton-symmetric configuration $(L_{e},L_{\mu},L_{\tau})=(0,0,0)$. For reference, the point of highest likelihood in this limited region of the parameter space is actually preferred to the “standard model” origin, with $2\ln(\tilde{\mathcal{L}}_{\mathrm{max}}/\tilde{\mathcal{L}}_{0})\approx-0.94$. For visualization purposes, we represent the confidence levels of the spectroscopic measurements of primordial abundances on the left panels of Fig. 4. The shift between the preferred regions for $Y_{p}$ and $\mathrm{D/H}$ stems from the “deuterium tension” in PRIMAT [Pitrou:2020etk], which will not be resolved or confirmed until better determinations of the deuterium burning nuclear rates are available [Pitrou:2021vqr].
@@
Figure 5 shows the evolution of neutrino asymmetries below $T=20\,\mathrm{MeV}$ for two configurations: on the top panel, a preferred point from the parameter space explored in Fig. 4 (shown with a black star), and on the bottom panel a configuration outside of the range of Fig. 4. Since the temperature range shown on Fig. 5 covers neutrino decoupling and electron-positron annihilations, we need to distinguish between the photon temperature $T_{\gamma}$ and the comoving temperature $T_{\mathrm{cm}}\propto a^{-1}$ (with $a$ the scale factor), which would be the neutrino temperature in the instantaneous neutrino decoupling approximation. Because of the variation of $T_{\gamma}/T_{\mathrm{cm}}$, the lepton number $L_{\alpha}$ defined in Eq. (3) is not a comoving quantity, which is why we multiply it by $(T_{\gamma}/T_{\mathrm{cm}})^{3}$ in Fig. 5 (see also the discussion in Ref. [Froustey:2024mgf]).
@@
These two examples result in a very small electron neutrino asymmetry at the BBN epoch, which is the main parameter determining BBN abundances. We note that future Simons Observatory and CMB-S4-like observations will be able to constrain $N_{\mathrm{eff}}$ at much better levels, and measure $Y_{p}$ at a level of precision comparable to current spectroscopic measurements [SimonsObservatory:2018koc, SimonsObservatory:2025wwn, CMB-S4:2016ple], thus being able to identify or constrain these required $L$ for the SF mechanism.
@@
With the results of Figs. 4 and 5, we highlight the variety of currently allowed asymmetry configurations for a given muon lepton number $L_{\mu}$. In the following, we then consider the lepton number $L$ which enters dark matter production as being unconstrained at the order $L\lesssim 10$ (recalling that it decreases by a factor of at least 5 between $T=10\rm\,GeV$ and $T=20\,\mathrm{MeV}$).
@@ II.3 Cosmological structure
[…]
We make the assumption that all lepton asymmetries are small by the time of structure formation. That is, they can correspond to lepton number configurations such as in the top panel of Fig. 5, so that the asymmetries are small enough that the structure formation is mostly sensitive to the sterile neutrino dark matter component, and not the small remaining active asymmetries. However, in general, SF mechanisms will likely be connected with nontrivial signals in $L_{\alpha}$ and $N_{\mathrm{eff}}$ in the BBN and CMB eras.
@@
and an example is shown in Fig. 6. We fit the SF transfer functions with thermal WDM transfer functions, where the thermal WDM particle mass is the free parameter for each sterile neutrino particle mass and mixing angle point. As shown, the residuals between the closest-fit thermal WDM transfer functions and those of the SF-produced sterile neutrinos are relatively small ($\lesssim 20\%$) for the suppression scale of 1/2. The $\lesssim 20\%$ accuracy holds in our fit across the examined parameter space, despite many points having PSDs that are highly nonthermal. In order to properly convert most of the prior literature’s inferred thermal WDM particle masses, $m_{\mathrm{th,old}}$ [Viel:2005qj] (more accurate at $m_{\mathrm{th}}\equiv m_{\mathrm{th,new}}\lesssim 5\,\mathrm{keV}$) into that calculated for larger thermal WDM particle masses, $m_{\mathrm{th,new}}$, found in Ref. [Vogel:2022odl], we use a linear fit relation,
@@
where the particle masses are in units of keV. This relation should be used to properly map any work which utilizes the previous WDM transfer function fit, in order to convert it into a more broadly calibrated thermal WDM particle mass value. We repeat our fitting process across the full parameter space and derive contours to describe the thermal WDM equivalent masses in order to apply signals and constraints directly to the SF parameter space. This is shown in Sec. III and Fig. 7.
Figure 7: Shown is the full sterile neutrino dark matter parameter space under SF production, with the initial $L$ values given at $T=10\rm\,GeV$ in order to achieve the observed density of the dark matter. These $L$ values are all reduced by a factor of 5 to 10 by $T=10\,\rm MeV$ from dark matter production and dilution (see Sec. II.2.1 and Fig. 3). The $L$ is further reduced by synchronized oscillations prior to BBN, and the initial $L$ values shown are consistent with BBN and CMB observations today (see Sec. II.2.2). The blue dot-dashed line is the best fit the thermal WDM particle mass from Villasenor+ [Villasenor:2022aiy]. The full region shown is consistent with the $1\sigma$ thermal WDM particle mass region from Villasenor+. The dashed red contour is the thermal WDM particle mass lower bound from the Zelko+ [Zelko:2022tgf], disfavoring the region left of the line. The gray cross-hatched x-ray exclusion region is the combination of NuSTAR and SPI/INTEGRAL at 95% CL upper limits [Perez:2016tcq, Roach:2019ctw, Roach:2022lgo, Krivonos:2024yvm, Fischer:2022pse].